OpenStudy (anonymous):

Please show all steps for the following word problem The lenght of rectangular yard is 12 feet more than twice its width. If the perimeter of the yard is 276, find the width

OpenStudy (anonymous):

x= width 2x+12= length You know the perimeter is 2 times the length plus two times the width. 2(x) + 2(2x+12)= 276 2x + 4x + 24= 276 6x + 24= 276 6x + 24 = 276 - 24 -24 6x= 252 x=42